OpenStudy (anonymous):

Solve the following problem for the roots by using the quadratic formula. 2(6 - x) = x(x + 5)

OpenStudy (anonymous):

1) Solve the following problem for the roots by completing the square. 2(6 - x) = x(x + 5)? here's your answer: Distributive property==> 6-2x=x squared+5x. Isolate variables==> 6=x sq+7x. Get everything onto on side==> 0=xsq+7x-6 (we subtracted 6 to get everything onto one side) Quadratic formula (see the link)==> http://2.bp.blogspot.com/_V8KsSIiGjBk/SsACMEj73KI/AAAAAAAAFIU/vNtErLdchMw/s1600/Quadratic%2BFormula.gif Substitute values and solve==> answers= -6 and 1. 2) Perform the indicated operation. (-3 + 5i) - (7 + 9i) Answer: Simplify (using order of operations)==> -3+5i-7-9i==> -4i-10. 3) Use the discriminant to determine the nature of the roots of the following equation. x2 + 2x + 5 = 0 Double root real and rational root real and irrational root imaginary root Answer: Discriminant= -18 (which means the equation only has 1 solution) Double root: None Real/rational root: None real/irrational root: none Imaginary root: Yes, we have imaginary roots 4) Last up: -10 + 4i 4 - 4i -10 - 4i Answer: leave them as they are. you can't simplify them anymore. Hope this helps! Source(s): http://2.bp.blogspot.com/_V8KsSIiGjBk/SsACMEj73KI/AAAAAAAAFIU/vNtErLdchMw/s1600/Quadratic%2BFormula.gif

OpenStudy (anonymous):

it says there is a square root in the eqaution

OpenStudy (anonymous):



OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i need to have one of those in it somehow

OpenStudy (anonymous):

@Luckyness

OpenStudy (anonymous):

hmm you have to go with the answer whats best for you